If someone takes a breath and the lungs expand from 4.3 L to 5.5 L in volume, and the initial pressure was 754 mm Hg, what is the pressure (in mm Hg) inside the lungs before any additional air is pulled in?
1
Expert's answer
2020-12-06T13:56:39-0500

4.3L to 5.5L


5.5L- 4.3L = 1.2L


initial pressure was 754 mm Hg,


P1 V1 = P2 V2


(754mmHg)(4.3L) = (P2)(5.5L)


3242.2 = 5.5 P2


P2 = (3242.2/5.5)


P2 = 589.49mmHg